a restaurant sells hamburgers and french fries. one family walks in a buys 4 hamburgers and 3 orders of fries for $10.23. A second family orders 1 hamburger and 5 orders of fries for $8.04. Write and slove a system of linear equations in order to find the cost of a hamburger and order of fries
Answer by mananth(16946)

x= cost of Hamburgers
y= cost of fries
Family I
4 x + 3 y = 10.23 .............1
Family II
1 x + 5 y = 8.04 .............2
Eliminate y
multiply (1)by -5
Multiply (2) by 3
-20 x -15 y = -51.15
3 x + 15 y = 24.12
Add the two equations
-17 x = -27.03
/ -17
x = 2
plug value of x in (1)
4 x + 3 y = 10.23
6 + 3 y = 10.23
3 y = 10.23 -6
3 y = 3.87
y = 1.29
x= 1.59 cost of Hamburgers
y= 1.29 cost of fries
m.ananth@hotmail.ca